For a transmission line the per unit length parameters are 0.1 Ohm/m, 0.2 uH/m, 10pF/m, and 0.02 Mho/m find the complex propagation constant at a frequency 1MHz

1
Expert's answer
2021-11-21T17:24:41-0500

"\gamma=\sqrt{ZY},"

"Z=R+i\omega L=0.1+0.2i,"

"Y=G+i\omega C=0.02+0.01i,"

"ZY=0.005i,"

"\gamma=±(0.05+0.05i)."